OpenStudy (anonymous):

PLEASE HELP WILL FAN AND GIVE MEDAL the sum of two consecutive odd integers is 244. what is the smaller integer. A. 119 B. 125 C. 123 D. 121

OpenStudy (zarkon):

\[\frac{244}{2}-1\]

OpenStudy (zarkon):

if the sum of two consecutive odd numbers was y then let x be the smaller number then x+(x+2)=y then 2x+2=y 2x=y-2 so \[x=\frac{y-2}{2}=\frac{y}{2}-1\] \[x=\frac{y}{2}-1\]